# A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models

### v1

@@ B.3. Data Statistics
Appendix G G. Localising Mover Heads
To efficiently classify the numerous attention heads (384 in GPT-2 medium) within the model, we calculate the Positional Patching Difference (PPD) to construct four-quadrant analysis. This method involves intervening in the attention value weights $W_{V}$ and constructs a distribution layout of heads, providing a systematic classification for quadrant groups. The layout is composed of the PPD score ($|S[p]|-|S[m_{1}]+S[m_{2}]|$) on the x-axis and the patching score for all sequence positions ($\mathcal{S}$) on the y-axis.
Based on this PPD score quadrant layout analysis, we categorise heads according to their quadrant positions, revealing insights into their functional roles:
@@
The intervention of attention head values ($W_{V}$) is crucial because all-term corruption maintains the same positional alignment for all samples $(s,m_{1},m_{2},p)\rightarrow(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$, resulting in minimal impact from attention head pattern interventions composed by attention query and key. Consequently, we can infer that the attention head output patching results primarily derive from the attention value activation. This property enables us to localise the source token positions from which the term information is moved by attention value patching.
Formally, we denote the attention query, key, value, and output matrix activations as $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$, and $\mathbf{W}_{O}$ respectively. One attention head output can be expressed as $\sigma\left((\mathbf{R}\mathbf{W}_{Q})(\mathbf{R}\mathbf{W}_{K})^{T}\right)(\mathbf{R}\mathbf{W}_{V})\mathbf{W}_{O}$, where $\sigma$ represents the non-linearity function (softmax in GPT-2) applied in the attention pattern and $\mathbf{R}$ denotes the residual stream.
@@
Our analysis identifies 9 heads for the symbolic circuit, as illustrated by red dots in Figure 8(a). $\mathbf{h}_{14.14}$, $\mathbf{h}_{15.14}$, and $\mathbf{h}_{18.12}$ are classified as positive copy heads, while $\mathbf{h}_{19.1}$ is grouped as a positive suppression head. $\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$, and $\mathbf{h}_{23.10}$ are classified as negative copy heads. It is noteworthy that non-symbolic results reveal a more complex and noisy pattern in the mover localisation process (Figure 8(b)).
Figure 8: Distribution of attention heads in PPD-based quadrant analysis. (a) Results for symbolic inputs. (b) Results for non-symbolic inputs. The x-axis represents normalised PPD scores, while the y-axis represents normalised attention value-based patching scores ($S$). Red dots indicate attention heads above the threshold ($\tau$) of patching score.
Appendix H H. Non-symbolic Circuit Analysis
@@
Appendix I I. Generalisability to Other Syllogisms
To assess the generalisability of our model, we applied the circuit ablation method to all 15 unconditionally valid syllogisms (Table 3) and the result is shown in Figure 10. Unconditionally valid syllogisms are logical arguments that maintain their validity irrespective of the truth values of their premises. The validity of these syllogisms is determined solely by their logical form, which is characterised by mood and figure combinations. The mood of a syllogism is defined by the arrangement of four proposition types (”All (A)”, ”No (E)”, ”Some (I)”, ”Some … not (O)”) across its two premises and conclusion. Conversely, the figure of a categorical syllogism is determined by the structural arrangement of terms within the constituent sentences. This approach enables the evaluation of syntactic logic in isolation from contextual interference and belief-bias effects.
@@
Appendix J J. Generalisability Across Model Sizes
To assess the generalisability of our findings across different model capacities, we extended our symbolic circuit analysis to various GPT-2 model sizes. Figure 11 presents the results of this comparative analysis.

### v2

@@ Appendix G Path Patching Details
Appendix H Localizing Mover Heads
To efficiently classify the numerous attention heads (384 in GPT-2 medium) within the model, we calculate the Positional Patching Difference (PPD) to construct a four-quadrant analysis:
$\text{PPD}=|S[p]|-|S[m_{1}]+S[m_{2}]|$
This method involves intervening in the attention value weights $W_{V}$ and constructs a distribution layout of heads, providing a systematic classification for quadrant groups. The layout is composed of the PPD score on the x-axis and the patching score for all sequence positions ($\mathcal{S}$) on the y-axis.
Based on this PPD score quadrant layout analysis, we categorize heads according to their quadrant positions, revealing insights into their functional roles:
@@ Appendix H Localizing Mover Heads
The intervention of attention head values is selected because all-term corruption maintains the same positional alignment for all samples $(s,m_{1},m_{2},p)\rightarrow(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$, resulting in minimal impact from attention head pattern interventions composed by attention query and key. Consequently, we can infer that the attention head output patching results primarily are derived from the attention value activation. This property enables us to localize the source token positions from which the term information is moved by attention value patching.
Formally, when we denote the attention query, key, value, and output weight matrices as $W_{Q},W_{K},W_{V}$, and $W_{O}$ respectively, and the residual stream vector as $r$, one attention head output can be expressed as: $\sigma\left((rW_{Q})(rW_{K})^{T}\right)(rW_{V})W_{O}$, where $\sigma$ represents the non-linearity function (softmax in GPT-2) applied in the attention pattern.
@@
Our analysis identifies 9 heads for the symbolic circuit, as illustrated by red dots in Figure 8(a). $\mathbf{h}_{14.14}$, $\mathbf{h}_{15.14}$, and $\mathbf{h}_{18.12}$ are classified as positive copy candidates, while $\mathbf{h}_{19.1}$ is grouped as a positive suppression candidates. $\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$, and $\mathbf{h}_{23.10}$ are classified as negative copy candidates. It is noteworthy that non-symbolic results reveal a more complex and noisy pattern in the mover localization process (Figure 8(b)).
Figure 8: Distribution of attention heads in PPD-based quadrant analysis. (a) Results for symbolic inputs. (b) Results for non-symbolic inputs. The x-axis represents normalized PPD scores, while the y-axis represents normalized attention value-based patching scores ($S$). Red dots indicate attention heads above the threshold ($\tau$) of the patching score.
Appendix I Non-symbolic Circuit Analysis
@@
Appendix J Generalizability to Other Syllogisms
To assess the generalizability of our model, we applied the circuit ablation method to all 15 unconditionally valid syllogisms (Table 3), and the result is shown in Figure 10. Unconditionally valid syllogisms are logical arguments that maintain their validity irrespective of the truth values of their premises. The validity of these syllogisms is determined solely by their logical form, which is characterised by mood and figure combinations. The mood of a syllogism is defined by the arrangement of four proposition types ("All (A)", "No (E)", "Some (I)", "Some … not (O)") across its two premises and conclusion. Conversely, the figure of a categorical syllogism is determined by the structural arrangement of terms within the constituent sentences. This approach enables the evaluation of syntactic logic in isolation from contextual interference and belief-bias effects.
@@
Model S BC BI
Group 1: Erratic Pattern
Pythia-70M 0.19 0.02 0.02
Pythia-160M 0.00 0.12 0.21
Group 2: Compatible Pattern
Pythia-410M 0.99 0.40 0.51
Pythia-1B 0.30 0.31 0.43
Llama-3.2-1B 1.00 0.81 0.70
Group 3: Variant Pattern
Qwen2.5-0.5B 1.00 0.96 0.90
Qwen2.5-1.5B 1.00 0.99 0.90
Table 5: Accuracy comparison of different models across three evaluation categories: Symbolic (S), Belief-Consistent (BC), and Belief-Inconsistent (BI). Group 1 models exhibit erratic activation patterns, failing to establish a stable reasoning mechanism. Group 2 models demonstrate a compatible reasoning circuit with consistent information flow and suppression. Group 3, particularly the Qwen models, exhibit a functional but distinct suppression mechanism, where suppression is applied at the final token rather than the $[p]$ position.
Appendix K Generalizability Across Model Sizes
Our approach is guided by the universality hypothesis, which suggests that similar representational patterns and circuits emerge across models trained on the same dataset (Olah et al., 2020; Kornblith et al., 2019; Li et al., 2015). We assume that core mechanisms should persist across these different model sizes. Figure 11 presents the results of this comparative analysis.
@@
Appendix L Generalizability Across Different Models
Figures 12 and 13 represent the results of applying the same circuit analysis in different model families. The $m$-suppression heads for each model were manually selected based on their likelihood, determined by patching scores and spatial positioning within the activation pattern. The corresponding model accuracies in the three datasets (Symbolic, Belief-Consistent, and Belief-Inconsistent) are presented in Table 5.
Figure 12: Comprehensive results of the symbolic circuit analysis across different series of models (Pythia). (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
Figure 13: Comprehensive results of the symbolic circuit analysis across different series of models (Qwen2.5, Llama3.2 and fine-tuned GPT-2 medium). (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
Appendix M Reproducibility Statement
Our study is conducted using TransformerLens (Nanda and Bloom, 2022) on an Nvidia A100 GPU with 80GB of memory. The dataset and code to reproduce our experiments are available online to encourage future work in the field [^1] http://anonynous_url.com.
Appendix N Statistical Significance
For all experiments, we use 90 samples each for both symbolic and non-symbolic arguments to balance analytical depth with computational efficiency. This sample size sufficiently yields statistically significant activation patching results ($p<0.05$).